OpenStudy (anonymous):

Find the perimeter of an object which has a length of square root of 75 feet and a width of square root of 48 feet. Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. What type of object might this be?

OpenStudy (campbell_st):

simplify the surds.. \[\sqrt{75} = \sqrt{25} \times \sqrt{3}\] and \[\sqrt{48} = \sqrt{16} \times \sqrt{3}\] then P = 2l + 2w it should be quite easy as all meaurements are multiples of root 3